OpenStudy (anonymous):

Give a recursive definition for the set X of all natural numbers that are one or two more than a multiple of 10. In other words, give a recursive definition for the set {1, 2, 11, 12, 21, 22, 31, 32, ... }. B1. 1 is in X. B2. ? is in X. R. If x is in X, so is x + 10

OpenStudy (amistre64):

split it into 2 parts .... even and odds

OpenStudy (amistre64):

for some n=1,2,3,4,\[a_{2n}=1+10(n-1)\]\[a_{2n+1}=2+10(n-1)\]

OpenStudy (amistre64):

might have to rewrite those 2n and 2n+1 ... i was thinking of n=0,1,2,3,4, ....

OpenStudy (anonymous):

Thanks so much, I see it now that I split it, I was able to solve B1 and R, now I see B2. 2 is in X

OpenStudy (amistre64):

\[a_{2n-1}=a_1+10(n-1)\]\[a_{2n}=a_2+10(n-1)\]\[a_1=1,~a_2=2\]

OpenStudy (anonymous):

thanks again for explain it

OpenStudy (amistre64):

youre welcome